A NEW MEASUREMENT OF THE ELECTRON’S ELECTRIC DIPOLE MOMENT USING YbF MOLECULES

J. J. HUDSON, D. M. KARA, I. J. SMALLMAN, B. E. SAUER, \underline{M. R. TARBUTT} and E. A. HINDS, Centre for Cold Matter, Blackett Laboratory, Imperial College London, Prince Consort Road, London SW7 2AZ, UK.

Using a beam of YbF molecules, we have made an improved measurement of the electron’s electric dipole moment. We present our new result and discuss how we plan to reduce the statistical and systematic uncertainties in future measurements.